# A Decade of Wheat Mapping for Lebanon

### v1

@@ Abstract
Crop Monitoring, Field Deliniation, Wheat Mapping, Parameter-Efficient Fine Tuning, Temporal Spatial Vision Transformer, Fields of The World
@@ I Introduction
  • Accurate Parcel-level Wheat Maping: Use precise field delineation to improve wheat mapping results.
  • Refined Post-Processing: Set of post-processing steps to improve field delineation results.
@@
Figure 1 shows that the porposed pipeline overcomes the stated challenges, allowing more reliable agricultural monitoring. An interactive version of this Figure can be accessed here: https://shorturl.at/PONJH.
@@ IV Performance Evaluation
In Figure  4, We conducted several experiments to fine-tune the different threshold values used in the proposed piepline.
@@ V Discussion and Future Work
To understand the spatial and temporal dynamics of wheat cultivation over the past decade, we analyzed changes in wheat areas, focusing on expansion, abandonment, and land-use transitions. Table II illustrates the instability of Lebanon’s economic and social landscape. Wheat cultivation in Lebanon experienced considerable fluctuations between 2016 and 2024, with peaks in 2019 and 2023 (same for number of fields), and significant drops in 2020 and 2024. The sharp decline in wheat cultivation from 2019 to 2020 reflects the impact of global events, such as the COVID-19 pandemic. More specifically, this substantial decrease between 2019 and 2020 strongly suggests the impact of the severe financial crisis and economic meltdown that hit the country at that time. Again in 2024, political turmoil and armed conflicts left a major impact on the wheat area. In addition to the numerical analysis presented in Table II, we use a Sankey diagram, a visual method to represent flows between different states to understand crop rotation practices.
@@
Figure 5-(a) shows a significant peak in 2019 (indicated by the thickest flow), suggesting strong incentives or favorable conditions for new wheat areas. The lowest annual new wheat area occurred in 2024, which corresponds to the armed conflict and labor availability.

### v2

@@ Abstract
Crop Monitoring, Field Delineation, Wheat Mapping, Parameter-Efficient Fine Tuning, Temporal Spatial Vision Transformer, Fields of The World
@@ I Introduction
  • Accurate Parcel-level Wheat Mapping: Use precise field delineation to improve wheat mapping results.
  • Refined Post-Processing: A set of post-processing steps to improve field delineation results.
@@
Figure 1 shows that the proposed pipeline overcomes the stated challenges, allowing more reliable agricultural monitoring. An interactive version of this Figure can be accessed here: https://shorturl.at/PONJH.
@@ IV Performance Evaluation
In Figure  4, We conducted several experiments to fine-tune the different threshold values used in the proposed pipeline.
@@ V Discussion and Future Work
To understand the spatial and temporal dynamics of wheat cultivation over the past decade, we analyzed changes in wheat areas, focusing on expansion, abandonment, and land-use transitions. Table II illustrates the instability of Lebanon’s economic and social landscape. Wheat cultivation in Lebanon experienced considerable fluctuations between 2016 and 2024, with peaks in 2019 and 2023 (same for number of fields), and significant drops in 2020 and 2024. The sharp decline in wheat cultivation from 2019 to 2020 reflects the impact of global events, such as the COVID-19 pandemic. More specifically, this substantial decrease between 2019 and 2020 strongly suggests the impact of the severe financial crisis and economic meltdown that hit the country at that time. Again in 2024, political turmoil and armed conflicts left a major impact on the wheat area. The resulting wheat maps can be publicly accessed at https://geogroup.ai/catalogue/#wheat
@@
In addition to the numerical analysis presented in Table II, we use a Sankey diagram, a visual method to represent flows between different states to understand crop rotation practices. Figure 5-(a) shows a significant peak in 2019 (indicated by the thickest flow), suggesting strong incentives or favorable conditions for new wheat areas. The lowest annual new wheat area occurred in 2024, which corresponds to armed conflict and labor availability.